OpenStudy (anonymous):

Solve for x. 10x + 12 = 132

Nnesha (nnesha):

12 is positive on left side in order to move 12 to the right side you have to do opposite so what is opposite of add ?

OpenStudy (anonymous):

Subtract.

Nnesha (nnesha):

yes so subtract 12 both side

OpenStudy (anonymous):

x = 12.

Nnesha (nnesha):

\(\huge\color{Green}{{\rm welcome}\rm~to~open~study!!!!}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{orange}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\) @B0og3r and @inowaIst both of ou welcome ! hope you guys read this http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

\[10x +12-12=13 2-12\]

OpenStudy (alexandervonhumboldt2):

yeah

Nnesha (nnesha):

yes right now 10 is multiplying by x what is opposite of multiply ?

OpenStudy (alexandervonhumboldt2):

so on left side 12-12=0 so you have 10x=120

OpenStudy (alexandervonhumboldt2):

now you have to divide both sides by 10 or just remove 10 from left side and on right side divide 120 by 10 and get the answer

OpenStudy (anonymous):

Opposite is divide.

Nnesha (nnesha):

yes right did you combine like terms first ??

Nnesha (nnesha):

10x+12−12=132−12 this one btw u already got that ^^^^ so u have 10x =120 opposite of multiply is divide so divide both side by 10

Nnesha (nnesha):

refresh page if there are soooo many question marks !

OpenStudy (anonymous):

I got 12.

Nnesha (nnesha):

okay to check your answer substitute 12 into the x and see if you get both side equal then that is your answer  both side equal check your answer is right :) good job